### v1

# Scaling Behaviors of LLM Reinforcement Learning Post-Training: An Empirical Study in Mathematical Reasoning

Abstract
While scaling laws for large language models (LLMs) during pre-training have been extensively studied, their behavior under reinforcement learning (RL) post-training remains largely unexplored. This paper presents a systematic empirical investigation of scaling behaviors in RL-based post-training, with a particular focus on mathematical reasoning. Based on 54 experiments across diverse model sizes and training settings, we characterize how model scale, data volume, and computational budget interact to shape performance. Our analysis leads to four key findings: ❶ Under a fixed computational budget, larger models trained for fewer steps consistently outperform smaller models trained for more steps. ❷ Given a fixed amount of training data, larger models achieve superior sample efficiency, yielding lower loss. ❸ In data-constrained regimes, repeated reuse of high-quality data proves highly effective, as final performance is primarily governed by the total number of optimization steps rather than the uniqueness of samples. ❹ These scaling behaviors are robust across both base and instruction-tuned models, which share similar learning dynamics (e.g., larger models show faster convergence) even while differing in absolute accuracy. Collectively, these results provide a principled foundation and practical guidelines for efficiently scaling the reasoning capabilities of LLMs through RL post-training.
1 Introduction
The rapid progress of large language models (LLMs) has made elucidating their scaling laws a matter of central importance. These laws, which capture the intricate relationships between model architecture, parameter size, computational cost, data availability, and downstream performance (Kaplan et al., 2020; Hoffmann et al., 2022), are invaluable not only because they illuminate the latent factors governing learning dynamics, but also because they provide actionable guidance on how to distribute scarce computational resources most effectively (Li et al., 2025a). While extensive efforts have clarified scaling behavior, the scaling behavior of reinforcement learning (RL) post-training for LLM reasoning remains underexplored.
During pretraining, Kaplan et al. (2020) show that cross-entropy loss follows smooth power-law scaling in model size, dataset size, and training compute, implying that larger models trained for fewer steps are compute-optimal. Hoffmann et al. (2022) refine this by showing that, under fixed compute, scaling parameters and tokens proportionally is optimal, since many large models are undertrained. Extending to neural-based RL, Hilton et al. (2023) empirically demonstrates that the intrinsic performance of convolutional neural networks (CNNs) optimized via reinforcement learning also scales power-law with model capacity and environment interaction.
Recently, RL has become the predominant post-training strategy for enhancing the reasoning abilities of LLMs, particularly in mathematics, a domain that demands long-horizon, compositional reasoning (Ferrag et al., 2025; DeepSeek-AI, 2025; Kimi Team, 2025; Ahn et al., 2024). Given the recent works applying RL for LLM reasoning, it is crucial to understand how to scale RL training effectively for reasoning tasks. In this work, we aim to fill this research gap with an extensive empirical study on models of varying scales. Defining the normalized test error rate as Test Loss ($L$), we formalize the resource allocation challenge as a series of constrained optimization problems:
  1. 1.
    The Compute-Constrained Scenario: Given a fixed computational budget $C$, we seek the model size $N$ (and corresponding data allocation $D$) that minimizes the final loss:
    $\underset{N,~D}{\operatorname*{arg\,min}}\;L(N,D)\quad\text{s.t.}\quad\mathrm{FLOPs}(N,D)=C_{\text{const}},$ (1)
  2. 2.
    The Data-Constrained Scenario: For a fixed amount of unique training data $D$, we aim to determine the model size $N$ that yields the lowest test loss under this data limitation:
    $\underset{N,~C}{\operatorname*{arg\,min}}\;L(N,C)\quad\text{s.t.}\quad D=D_{\text{const}},$ (2)
  3. 3.
    The Data Reuse Scenario: With a fixed number of optimization steps $S$, we investigate the impact of the data reuse factor $\tau$—the number of times each unique example is revisited—on the final loss. In this setting, the total number of processed samples is constrained as $D_{\text{unique}}\times\tau=S_{\text{const}}$. Formally:
    $\underset{\tau,~C}{\operatorname*{arg\,min}}\;L(\tau,C)\quad\text{s.t.}\quad D_{\text{unique}}\times\tau=S_{\text{const}},$ (3)
To empirically answer these questions, we fine-tune 54 LLMs with reinforcement learning on 50k mathematics problems, based on the Qwen2.5 model family (Qwen et al., 2025). Figure 1 shows that, within the 0.5-14B range, and tested in an environment close to the training set distribution, the loss reduction brought by RL follows an approximately log-linear trend with compute and data size. Importantly, larger models not only have better initial performance but also generally have more efficient computation and data utilization during the optimization process.
(a) Relation between test loss and training compute for base models with fitted curve.
(b) Relation between test loss and data size for base models with fitted curve.
Figure 1: Key finding of this study: In both cases, larger models consistently outperform smaller models at all points in training.
We further analyze the data-constrained regime, where we demonstrate that data reuse is a highly effective strategy. We validate the generality of our findings through extensive ablation studies on both base and instruct model series. Besides, we also study the impact of the key hyperparameter rollout number in the GRPO algorithm (Shao et al., 2024). These investigations establish fundamental scaling relationships for RL post-training, providing a quantitative foundation and practical guidelines for resource-efficient model refinement.
Specifically, our key findings can be summarized as follows:
  • In our experiment scale, larger models, starting with stronger initial performance, consistently achieve better compute and data efficiency in RL post-training for mathematical reasoning.
  • In data-limited settings, repeated exposure to a small dataset is nearly as effective as using larger corpora, highlighting data reuse as a practical strategy.
  • In math reasoning, scaling behaviors are robust across both base and instruction models and GRPO hyperparameters, providing actionable guidance for resource-efficient RL post-training.
2 Experimental Setup
We describe the experimental setup for studying scaling behavior in RL post-training of LLMs for mathematical reasoning, including the model family, training and evaluation data, and evaluation protocol in this section. Full details are provided in Appendix A.
Models and Framework.
We use the Qwen2.5 model family (0.5B, 1.5B, 3B, 7B, and 14B parameters) (Qwen et al., 2025), which shares the same architecture, so that parameter count is the only variable in our scaling analysis. All experiments are run with the VeRL framework (Sheng et al., 2024), a large-scale RL platform for LLMs ensuring consistency and reproducibility.
RL Algorithm.
[…]
$\mathcal{L}_{\text{GRPO}}=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\lvert o_{i}\rvert}\sum_{t=1}^{\lvert o_{i}\rvert}\Bigl\{\min\Bigl[\rho(\theta)\,\hat{A}_{i,t},\;\operatorname{clip}\!\Bigl(\rho(\theta),1-\varepsilon,\,1+\varepsilon\Bigr)\hat{A}_{i,t}\Bigr]-\beta\,\mathrm{D}_{\mathrm{KL}}\Bigr\}.$ (4)
Where $\rho(\theta)=\frac{\pi_{\theta}\!\left(o_{i,t}\mid q,\,o_{i,<t}\right)}{\pi_{\theta_{\mathrm{old}}}\!\left(o_{i,t}\mid q,\,o_{i,<t}\right)}$ is the important sampling weight. For each output $o_{i}$, a reward model or rule is used to yield the reward signal $\mathbf{r}=\{r_{1},r_{2},\cdots,r_{G}\}$. The advantage is computed as
$\hat{A}_{i,t}=\frac{r_{i}-\mathrm{mean}(\mathbf{r})}{\mathrm{std}(\mathbf{r})}.$ (5)
Dataset settings.
The training data is the mathematics subset of the guru-RL-92k dataset from the Reasoning360 project (Cheng et al., 2025), which is carefully curated through deduplication and difficulty filtering. We further sort the problems by increasing difficulty (decreasing pass rate) to enable curriculum learning. The evaluation data consists of two parts. To derive scaling laws, we use a held-out set of 500 in-domain math problems sampled from the training distribution. To assess generalization, we evaluate on a broader benchmark suite spanning mathematics (AIME2024 (Patel et al., 2024), AMC2023 (KnovelEng, 2025), GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023)), code (HumanEval (Chen et al., 2021)), logic (Zebra Puzzle(Lin, 2024)), and science (SuperGPQA (Team et al., 2025)). More details about dataset settings can be found in Appendix A.1.
Prompt Setting.
To ensure stable behavior during RL training and evaluation, we use structured prompts tailored to each domain. For example, all mathematics problems are prepended with the Chain-of-Thought prompt (Wei et al., 2023): “You are a knowledgeable math assistant. Answer the following questions and think step by step”. More prompt templates for all related domains could be found in Appendix A.3.
Evaluation Process.
[…]
Metric.
Our primary evaluation metric is the test loss ($L$), a proxy for reward-based performance in the RL setting. Unlike conventional RL, where agents interact with an environment and sample rewards freely, LLM post-training mainly focuses on the fixed data set and a verifiable reward rule. We hold out part of the training corpus and compute the normalized error rate. Formally, $L=1-(R/R_{\text{max}})$, where $R$ is the number of correct solutions and $R_{\text{max}}$ the total. Obviously, the goal of maximizing reward in RL training is equivalent to minimizing $L$.
3 Empirical Results and Scaling Laws
This section presents a comprehensive empirical investigation into the scaling behavior of RL for post-training LLMs. Our experiments are designed to address the core questions posed in Eq. 1, Eq. 2, and Eq. 3. We first examine scaling behaviors under compute and data constraints, then analyze independent scaling dimensions, data reuse strategies, and finally evaluate generalization performance together with an ablation study on the GRPO group size $G$ (Eq. 4). To ensure robust conclusions, each configuration is repeated three times for both base and instruct models.
3.1 Compute-Optimal Scaling
Observation 1 Within the model parameter range of 0.5B to 14B, RL post-training under a fixed computational budget $C$ is compute-optimal when prioritizing the training of larger models rather than extending the training duration of smaller models.
(a) Base Models
(b) Instruct Models
Figure 2: Test loss as a function of training compute ($C$). Larger models consistently achieve lower loss at the same compute budget, showing their higher computational efficiency.
In the compute-constrained setting (Eq. 1), we train 0.5B–14B models and measure test loss as a function of cumulative FLOPs $C$. As shown in Figure 2, larger models consistently outperform smaller ones under the same compute budget for both base and instruct variants. The loss–compute relationship follows a log-linear trend, which can be modeled by a power law:
$\log(L(N,C))=-k_{C}(N)\cdot\log(C)+E_{C}(N).$ (6)
Table 1 shows $k_{C}(N)$ increasing with model size $N$, indicating that larger models not only start from lower loss but also yield greater performance gains per FLOP. This contrasts with pre-training scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), which observe a compute-efficient frontier of the trade-off between model size and computation. In our experiments, no such trade-off is observed, which we believe is due to pre-training allocating substantially greater computational resources and stronger capabilities to larger models that subsequently dominate the post-training dynamics.
This result observed from tested models (0.5B-14B) suggests that, in RL post-training, it is generally preferable to allocate resources toward training larger models. Nevertheless, we hypothesize that this trend may not hold indefinitely, as increasingly larger models start from lower initial loss with making more challenging loss reduction.
Table 1: Fitted slopes $k_{C}$ for different model sizes.
Model 0.5B 1.5B 3B 7B 14B
Base 0.0075 0.0338 0.0168 0.0370 0.0680
Instruct 0.0055 0.0063 0.0153 0.0436 0.0454
3.2 Data-Optimal Scaling
Observation 2 Within the model range of 0.5B to 14B, for a fixed volume of unique training data $D$, larger models demonstrate superior sample efficiency, consistently achieving lower test loss.
(a) Base Models
(b) Instruct Models
Figure 3: Test loss as a function of data size $D$. Larger models consistently outperform smaller ones given the same number of unique samples.
In the data-constrained setting (Eq. 2), we train models with varying parameter counts on fixed amounts of unique samples. As shown in Figure 3, larger models consistently achieve lower test loss across both base and instruct variants. The loss–data relationship follows a power law:
$\log(L(N,D))=-k_{D}(N)\cdot\log(D)+E_{D}(N)$ (7)
with the sample efficiency coefficient $k_{D}(N)$ increasing with model size (Table 2). It confirms that larger models not only compute more efficiently, but also extract more knowledge from each data point, showing that scaling up model size is the most effective strategy for RL post-training. This observation is similar to pre-training behavior (Kaplan et al., 2020; Hoffmann et al., 2022).
Table 2: Fitted slopes $k_{D}$ for different model sizes.
Model 0.5B 1.5B 3B 7B 14B
Base 0.0070 0.0268 0.0181 0.0409 0.0739
Instruct 0.0054 0.0065 0.0164 0.0469 0.0504
3.3 Scaling up Model Size
Observation 3 When trained to converge on sufficiently large datasets, test loss decreases monotonically with model size, though the trend deviates from a strict power law.
We train models of varying sizes to convergence and compare their final test loss. As shown in Figure 4(a), larger models consistently achieve lower loss, improving monotonically with scale. The curve deviates from a strict power law: smaller models show weaker gains, suggesting diminishing returns at low parameter counts. A likely reason is that larger models inherit richer pre-trained representations, which reinforcement fine-tuning exploits for greater improvements than parameter growth alone would predict. Figure 4(b) further shows that as RL training progresses, larger models generate longer responses. This correlates with higher accuracy, indicating greater test-time scaling efficiency: additional inference tokens yield larger gains in bigger models.
We also benchmark our RL-tuned Qwen2.5 models (Qwen et al., 2025) against state-of-the-art open-source reasoning systems, including Qwen3 (Yang et al., 2025) and GPT-OSS (OpenAI et al., 2025). On our held-out set, the 7B and 14B models match or surpass dense Qwen3 counterparts of similar size, highlighting the effectiveness of RL post-training. Mixture-of-experts models such as Qwen3 and GPT-OSS achieve lower loss at much larger scales (20B–235B), with GPT-OSS-120B currently leading. These comparisons suggest that scaling beyond 14B will be necessary to fully characterize post-training behavior and compete with frontier MoE systems.
(a) Relation between test loss and model size NN when trained to convergence on sufficiently large data.
(b) Relation between response length and data size for instruct models.
Figure 4: Analysis of model scaling properties.
3.4 Scaling with Constrained Data and Reuse
Observation 4 Performance in data-constrained settings is primarily determined by the total number of parameter update steps ($S$). For a fixed $S$, the final test loss is remarkably insensitive to the data reuse factor ($\tau$), with no significant degradation up to $\tau=25$.
(a) Base Model
(b) Instruct Model
Figure 5: Plot shows the final test loss of a 7B Base (left) and Instruct (right) model trained with a fixed toal number of samples but varying the data reuse factor $\tau$.
We further consider the data-reuse scenario (Eq. 3), where high-quality data is limited but can be revisited multiple times. To simulate this, we partition the training set into smaller subsets while preserving the difficulty distribution (Details provided in Appendix A.4). Each subset is cycled through multiple times, with the reuse factor $\tau$ controlling how often each unique example is revisited. The total number of update steps $S$ is fixed across runs, and curriculum ordering is maintained so that problems are always presented from easy to hard. This ensures that performance differences arise solely from the degree of data reuse, rather than distributional or scheduling artifacts.
As shown in Figure 5, performance remains nearly unchanged for $\tau\leq 25$, while moderate degradation appears as $\tau$ increases further. At $\tau=100$, we observe clear signs of overfitting, indicating that repeated reuse eventually harms generalization. Overall, these results suggest that final performance is primarily governed by the total number of optimization steps rather than sample uniqueness, and that moderate data reuse is an effective strategy for RL fine-tuning with limited datasets.
3.5 Domain Transfer
Observation 5 RL post-training on mathematical reasoning yields generalization improvements on in-domain tasks with varying difficulty, but shows negligible transfer to out-of-domain tasks.
(a) A consistent decrease in test loss is observed on various in-domain mathematics tasks.
(b) Mixed results are observed on out-of-domain tasks.
Figure 6: The effect of domain transfer, illustrated with the Qwen2.5-14B-Base model.
We also investigate the generalization capabilities of the reinforcement learning fine-tuning (RFT) models by evaluating them on a suite of unseen in-domain tasks with varying difficulty and out-of-domain (OOD) tasks. More results are in Appendix B.1.
In-Domain Generalization. Figure 6(a) shows consistent improvements on unseen mathematics tasks outside the training set. On benchmarks, from easy to hard, including GSM8K, MATH-500, AMC2023, AIME2024, test loss steadily decreases with training compute, suggesting that RL post-training enhances transferable reasoning skills within mathematics.
Out-of-Domain Generalization. As shown in Figure 6(b), results on OOD tasks are markedly different. For code generation (HumanEval) and STEM problems (SuperGPQA), performance gains marginally, indicating that RL fine-tuning is highly specialized. More strikingly, on logical reasoning (zebra_puzzle), the 14B model’s test loss increases during training, suggesting that intensive domain-specific optimization can negatively impact distinct reasoning skills.
3.6 Ablation on GRPO Hyperparameters
Observation 6 A larger GRPO rollout group size ($G$) is consistently more data-efficient, while the compute-optimal $G$ grows with the total computational budget.
We conducted an ablation study on the rollout group size $G$, a key GRPO hyperparameter that controls how many responses are sampled per prompt. This directly affects both the compute per update and the stability of the training signal. We tested $G\in\{4,8,16,32\}$ on the 7B models.
Data-centric View. Figure 7(b) and 7(d) shows that larger rollout sizes consistently yield better sample efficiency: $G=32$ achieves the lowest test loss for the same number of unique samples. This supports the intuition that more responses per question provide a stronger advantage estimate and thus more effective gradient updates.
Compute-centric View. The optimal rollout size $G$ is not fixed but shifts with the training budget. This implies that practitioners should tune $G$ according to available compute rather than relying on a universal setting. We attribute this dynamic to the trade-off between the higher variance reduction from larger $G$ and the additional FLOPs it consumes, which makes small $G$ preferable at low budgets but large $G$ superior when ample compute is available.
(a) 7B-Base: Loss vs. Compute
(b) 7B-Base: Loss vs. Data Size
(c) 7B-Instruct: Loss vs. Compute
(d) 7B-Instruct: Loss vs. Data Size
Figure 7: Effects of GRPO rollout size on training efficiency: (a–b) show compute/data scaling for 7B-Base; (c–d) for 7B-Instruct.
4 related work
Foundational Scaling Laws of Neural Language Models. Foundational scaling studies show language modeling loss follows smooth power-laws in model size $N$, data $D$, and compute $C$ (Kaplan et al., 2020), with compute-optimal training prescribing near lockstep growth of parameters and tokens under fixed FLOPs (Hoffmann et al., 2022). Later analyses attribute earlier discrepancies to embedding/non-embedding parameter accounting, last-layer costs, optimizer warmup, and scale-sensitive hyperparameters (Pearce & Song, 2024; Porian et al., 2024), while data-centric refinements examine pruning efficiency (Sorscher et al., 2022), repetition effects (Hernandez et al., 2022), gzip-based complexity predictors (Pandey, 2024), constrained or synthetic regimes (Muennighoff et al., 2023; Qin et al., 2025), and task transfer (e.g., translation) (Isik et al., 2024). Test-time compute amplification supplies an inference analogue to classical training laws (Snell et al., 2024).
RL post-training in LLMs. In RL, power-law trends similarly link capacity, interaction compute, and performance (Hilton et al., 2023); scaling RFT across horizon and compute improves mathematical and coding reasoning (DeepSeek-AI, 2025; Kimi Team, 2025; Mai et al., 2025b; Zhang et al., 2025a; b), while extended schedules (Liu et al., 2025), ultra-low-shot or single-example RL (Wang et al., 2025), and minimal-data efficiency paradigms (Li et al., 2025b) probe data–compute tradeoffs. Instability and uneven gains highlight fragile optimization (Zeng et al., 2025a; Yue et al., 2025), and multi-domain mixtures reveal both synergy and interference across math, code, and logic (Li et al., 2025c; Cheng et al., 2025).
Mathematical Reasoning with LLMs. Mathematical reasoning amplifies these dynamics: accuracy generally scales upward while verification behaviors remain inconsistent (Touvron et al., 2023); corpus volume and quality jointly shape attainable curves (Ye et al., 2024); multi-task math-generalist training diverges from specialist scaling trajectories (Yue et al., 2023); and RL with code execution induces additional behaviors such as emergent tool use concentrated in math problem solving (Zeng et al., 2025b). Collectively, evidence indicates that reasoning performance is governed by interacting axes of model size, data distribution/quality, training (supervised vs. RL) paradigm, and allocation of both training and inference compute, while unified laws for mathematical reasoning remain only partially characterized.
5 Discussion
Scaling Dependence on Evaluation Environment and Metrics. Reinforcement learning optimizes directly for environment rewards (Sutton & Barto, 2018), which in principle allows unbounded capability—as demonstrated by AlphaZero mastering board games (Silver et al., 2017), AlphaFold predicting protein structures (Jumper et al., 2021), and frontier LLMs such as Gemini-2.5-Pro achieving IMO-level performance (Huang & Yang, 2025). In contrast, text-based LLMs lack a well-defined RL environment, forcing us to rely on human-curated datasets as proxies. Test loss thus serves as a pragmatic but imperfect metric: it is monotonic and convergent, yet heavily dependent on dataset construction and task difficulty, with different benchmarks (e.g., GSM8K vs. AIME, Section 3.5) showing distinct convergence rates. This task dependence makes the absolute coefficients of our fitted scaling laws ($k(N),E$) difficult to interpret universally. Prior work proposed “intrinsic performance”—the minimum compute needed to reach a target reward—as a normalization across environments (Hilton et al., 2023), but we did not find an analogous measure in large-scale LLMs. Establishing principled, environment-independent evaluation protocols remains an open and critical challenge for RL-based scaling studies.
Scaling Dependence on Model Scale. Our study of models from 0.5B to 14B parameters shows that larger models exhibit greater sample and compute efficiency in RL post-training. This parameter range, however, is still below that of frontier models. We anticipate these advantages may not extend indefinitely, as performance can saturate on simple benchmarks like GSM8K and fail to converge on extremely hard problems. A plausible hypothesis is the existence of a ”sweet spot” or critical model scale—large enough for efficient RL adaptation but not so over-parameterized that returns diminish. Identifying this optimal scale is still an important direction for future work.
Dependence on RL Algorithm. Our analysis is based on GRPO, a mainstream and stable RL post-training algorithm for LLMs that uses an actor-only design and normalizes rewards across responses. Comparative study with alternative RL algorithms  (Cui et al., 2025) reports minor differences in training curves. Whether more advanced algorithms can significantly improve sample efficiency or stability—and thereby reshape the scaling frontier—remains an important open question.
Future of LLM Agent. The integration of reinforcement learning with agentic LLMs is increasingly viewed as a promising direction (Zhang et al., 2025a; b). Both theoretical and empirical studies show that augmentations such as external tool use and long-term memory can substantially boost model performance (Lin & Xu, 2025; Houliston et al., 2025; Mai et al., 2025a). We anticipate that such agentic mechanisms will markedly improve the scaling behavior of RL-trained LLMs: by offloading deterministic computations to tools and focusing learning on high-level decision making, these models could achieve much higher efficiency, effectively shifting the performance frontier upward for a given compute or data budget. Understanding the scaling laws of these agentic systems is, therefore, a key and exciting avenue for future research.
6 Conclusion
This study presents the systematic exploration of scaling laws for reinforcement learning post-training of large language models in mathematical reasoning. Through 54 controlled experiments across the Qwen2.5 family, we show that larger models consistently achieve superior compute and data efficiency, that performance in data-limited regimes depends primarily on the total number of optimization steps rather than data uniqueness, and that moderate data reuse can be highly effective without harming generalization. While RL post-training reliably strengthens in-domain reasoning, its transfer to out-of-domain tasks remains limited, underscoring the trade-off between specialization and breadth. Our ablation further identifies rollout size in GRPO as a practical lever tied to compute budgets. Taken together, these findings offer principled and actionable guidelines for resource-efficient RL fine-tuning and suggest promising directions for further exploration of scaling and generalization in LLM reasoning.
7 Statement
7.1 Ethics statement
The authors have read and adhered to the ICLR Code of Ethics. This work is foundational in nature, focusing on the scaling properties of large language models in the domain of mathematical reasoning. Our research exclusively utilizes publicly available and previously published resources, including open-source models (e.g., Qwen2.5) and established datasets (e.g.,guru-RL-92k), thereby mitigating concerns related to data privacy, human subjects, or the release of sensitive information. The application domain of mathematical problem-solving does not inherently present risks of direct societal harm. The primary ethical consideration associated with this work is the environmental impact of the computational resources required for large-scale model training, a challenge common to the field. We believe that by providing insights into efficient resource allocation, our work contributes positively to mitigating this concern for future research.
7.2 Reproducibility statement
We are committed to ensuring the reproducibility of our research. All aspects of our experimental design are detailed to facilitate this. The specific models used (Qwen2.5 series), are described in Section 2. The training and evaluation datasets, including our held-out set, are also detailed in Section 2, with a full breakdown provided in Appendix A.1. Our reinforcement learning methodology is based on the GRPO algorithm, and all relevant hyperparameters used for training and evaluation are listed in Appendix A.2. The evaluation protocol, including the precise definition of our Test Loss metric and the prompt templates for all tasks, is outlined in Section 2 and Appendix A.3. All figures and observations presented are directly generated from the raw data collected during our experiments. At last, we provide the experiment source code in Supplementary Material along with the submission.
Appendix A Experiment Setup Details
This section provides a detailed breakdown of the datasets and hyperparameters used in our experiments, supplementing the information provided in the main text.
A.1 Dataset Details
Our training was conducted on a curated mathematics dataset. For evaluation, especially for analyzing generalization (as mentioned in the main text), we utilized a comprehensive suite of benchmarks spanning multiple domains. The composition of this evaluation suite is detailed in Table 3.
Table 3: Composition of the multi-domain evaluation suite.
Dataset Samples Huggingface Tag Domain
Held-out Data 500 LLM360/guru-RL-92k Math
aime2024 30 Maxwell-Jia/AIME_2024 Math
amc2023 40 knoveleng/AMC-23 Math
codegen_humaneval 164 openai/openai_humaneval Code
gsm8k 1319 openai/gsm8k Math
logic_zebra_puzzle 200 LLM360/guru-RL-92k Logical Reasoning
math 500 HuggingFaceH4/MATH-500 Math
stem_supergpqa 200 LLM360/guru-RL-92k STEM
Total 2953
A.2 Hyperparameter Configuration
All experiments were conducted with a consistent set of hyperparameters for the Group Relative Policy Optimization (GRPO) algorithm to ensure a fair comparison across different model sizes and configurations. The key hyperparameters are listed in Table 4.
Table 4: GRPO training hyperparameters used across all experiments.
Hyperparameter Value
Learning Rate $1.0\times 10^{-6}$
Batch Size 512
KL Loss Coefficient 0.001
Rollout Temperature (Training) 1.0
Rollout Temperature (Evaluation) 0.7
Clip Ratio (High & Low) 0.2
Input Sequence Length 2048
Output Sequence Length 4096
A.3 Prompt Templates
This section details the specific prompt templates used for evaluating models on different domains. For each task, the model was provided with the corresponding instruction prepended to the problem statement <question>.
Table 5: Prompt templates used for different evaluation domains.
Domain Prompt Template
[…]
A.4 Data Reuse Experiment Setup
Figure 8: Data Reuse Schema
To systematically evaluate the effect of data reuse under constrained data scenarios, we design controlled experiments where all runs are trained with the same total data size but different levels of data repetition. Each run randomly samples a subset from the full training corpus and repeats this subset sufficiently many times to exactly match the target data budget (i.e., $\text{subset size}\times\tau=\text{total data size}$). Unlike Muennighoff et al. (2023), subsets are sampled independently for each run rather than sampling within the larger subsets, to mitigate sampling bias and balance stochasticity across conditions. To remain consistent with the Curriculum Learning setting of the main experiments, examples within each subset are ordered by increasing difficulty; across epochs, this difficulty schedule is preserved and repeated rather than reshuffled, as illustrated in Figure 8.
Appendix B Additional Experiment Results
This section provides supplementary experimental results that support and extend the analyses presented in the main body of the paper.
B.1 Performance on In-Domain and Out-of-Domain Tasks
To assess how the mathematical reasoning capabilities acquired during RL fine-tuning generalize, we evaluated our models on a comprehensive suite of unseen benchmarks. We categorize these into two groups: in-domain different tasks (other mathematics datasets) and out-of-domain tasks (e.g., code, science, logic). The results are presented in Figure 9 and Figure 10.
Figure 9: Test loss on in-domain and out-of-domain benchmarks vs data size for Base models. It shows modest positive transfer on in-domain tasks, with limited or negative transfer on OOD tasks.
In-Domain Generalization (Different Mathematical Tasks). On mathematics benchmarks not included in our training set (such as GSM8K, MATH, AIME, and AMC), we observe a generally positive transfer of learned skills. For most of these tasks, the test loss shows a modest but consistent decrease as training progresses, particularly for the larger models. This suggests that the model’s enhanced reasoning ability is not overfitted to the training distribution and is applicable to a wider range of mathematical problems.
Out-of-Domain Generalization. When evaluating on tasks outside of mathematics, the generalization is more limited. For both code generation (HumanEval) and science problems (SuperGPQA), performance remains largely static throughout the training process across all model sizes, with test loss curves staying flat. This indicates that the specialized mathematical reasoning skills do not readily transfer to these domains. A noteworthy phenomenon is observed in the logical reasoning task (Zebra Puzzle): the largest models (particularly the 14B variants) show a degradation in performance (an increase in test loss) as training progresses, suggesting a potential negative transfer effect where intensive optimization on mathematical reasoning may interfere with capabilities required for certain types of logical puzzles.
Figure 10: Test loss on in-domain and out-of-domain benchmarks vs data size for Instruct models. It shows modest positive transfer on in-domain tasks, with limited or negative transfer on OOD tasks.
B.2 Peformance Compared with Advance Models
Model Family Model Identifier Pass@1 Score
Models from Our Study (Post-RFT)
Qwen2.5-Base 0.5B 0.070
1.5B 0.116
3B 0.182
7B 0.338
14B 0.450
Qwen2.5-Instruct 0.5B 0.078
1.5B 0.138
3B 0.216
7B 0.380
14B 0.488
External SOTA Models (for Comparison)
Qwen3 0.6B 0.178
1.7B 0.288
4B 0.418
8B 0.366
14B 0.388
30B (A3B) 0.528
32B 0.412
235B (A22B) 0.602
GPT-OSS 20B 0.556
120B 0.660
Table 6: Performance of various models on the held-out evaluation set.
To contextualize the performance of our models and the difficulty of our primary evaluation metric, we benchmarked a range of external, state-of-the-art (SOTA) models on our held-out mathematics test set. The results are presented in Table 6. The performance of our Qwen2.5 models reflects their final scores after the completion of reinforcement learning fine-tuning (RFT), while others are benchmarked directly.
Appendix C Formula Fitting and Derivation
C.1 FLOPs Calculation Methodology
The computational cost for a LLM is primarily determined by the number of non-embedding parameters ($N$) and the number of processed tokens ($T$). The costs for the fundamental operations are:
  • Forward Pass Cost: The cost of a single forward pass is approximately $C_{\text{fwd}}\approx 2NT$ FLOPs.
  • Backward Pass Cost: The backward pass is approximately twice as expensive as the forward pass, so $C_{\text{bwd}}\approx 4NT$ FLOPs.
A full training step, which includes one forward and one backward pass for the gradient update, therefore has a total computational cost of:
$C_{\text{train}}=C_{\text{fwd}}+C_{\text{bwd}}\approx 2NT+4NT=6NT\text{ FLOPs.}$ (8)
$\text{FLOPs}_{\text{step}}=6\times N\times T_{\text{step}}$ (9)
By recording the exact number of processed tokens $T$ per step, we compute the cumulative FLOPs reported throughout this paper as the sum of these per-step calculations over the course of training.
C.2 Coefficient Comparison
Figure 11: Fitted coefficients $k(N)$ and $E_{D}(N)$ comparison over $L(N,D)$ and $L(N,C)$. Upper: base model. Bottom: instruct model.
We consider the two laws
$\ln L(N,C)=-\,k_{C}(N)\,\ln C+E_{C}(N),$ (10)
and
$\ln L(N,D)=-\,k_{D}(N)\,\ln D+E_{D}(N),$ (11)
are consistent under the linkage $C=ND\phi$ where $\phi>0$ is a constant for simplification.
Claim. Under $C=ND\phi$, the slopes coincide and the intercepts differ by a known shift:
$k_{C}(N)$ $=k_{D}(N)\;=\;k(N),$ (12)
$E_{C}(N)$ $=E_{D}(N)\;+\;k(N)\,\ln\!\bigl(N\phi\bigr).$ (13)
Proof. Substitute $C=ND\phi$ into equation 10:
$\ln L(N,C)$ $=-\,k_{C}(N)\,\ln\!\bigl(ND\phi\bigr)+E_{C}(N)$
$=-\,k_{C}(N)\,\bigl[\ln D+\ln(N\phi)\bigr]+E_{C}(N)$
$=-\,k_{C}(N)\,\ln D\;+\;\Bigl(E_{C}(N)-k_{C}(N)\,\ln(N\phi)\Bigr).$
Comparing this with equation 11, i.e., $\ln L(N,D)=-\,k_{D}(N)\,\ln D+E_{D}(N)$, equality for all $D>0$ forces the coefficients of $\ln D$ and the constants to match:
$k_{D}(N)=k_{C}(N)=:k(N),\qquad E_{D}(N)=E_{C}(N)-k(N)\,\ln(N\phi).$
Rearranging the second identity yields equation 13. $\square$
The observation from Figure 11 also matches with this conclusion.
C.3 Fitting for K and E
Table 7: $L(N,C)$ Fitting Results: $\log\text{L}=-k_{C}\log C+E_{C}$
Model Size Base Instruct
$k_{C}$ $E_{C}$ $R_{C}^{2}$ $k_{C}$ $E_{C}$ $R_{C}^{2}$
0.5B 0.0075 0.1140 0.485 0.0055 0.0744 0.313
1.5B 0.0338 0.5889 0.813 0.0063 0.0718 0.293
3B 0.0168 0.2448 0.621 0.0153 0.2013 0.570
7B 0.0370 0.5562 0.789 0.0436 0.6600 0.809
14B 0.0680 1.1023 0.921 0.0454 0.6361 0.817
Table 8: $L(N,D)$ Fitting Results: $\log\text{TestLoss}=-k_{D}\log D+E_{D}$
Model Size Base Instruct
$k_{D}$ $E_{D}$ $R_{D}^{2}$ $k_{D}$ $E_{D}$ $R_{D}^{2}$
0.5B 0.0070 0.0117 0.483 0.0054 0.0002 0.314
1.5B 0.0268 0.0839 0.796 0.0065 -0.0150 0.294
3B 0.0181 0.0127 0.617 0.0164 -0.0112 0.573
7B 0.0409 0.0359 0.792 0.0469 0.0399 0.812
14B 0.0739 0.1200 0.921 0.0504 -0.0167 0.818
Appendix D Response Length
Figure 12: Response length vs. Data size. Left: Base models. Right: Instruct models.
Figure 13: Left: Fitted response length vs. data size (Instruct models). Right: The slopes $k_{response}(N)$
Appendix E Statements
E.1 The Use of Large Language Models
We used Large Language Model (LLM) to refine our initial draft. This process included checking for obvious grammatical and syntactical errors, as well as making the language more formal and academic. We reviewed the content generated by the LLM to ensure that no prohibited generated content appeared in the article.