### v3

# Scaling Behaviors of LLM Reinforcement Learning Post-Training: An Empirical Study in Mathematical Reasoning

Abstract
While scaling laws for large language models (LLMs) during pre-training have been extensively studied, their behavior under reinforcement learning (RL) post-training remains largely unexplored. This paper presents a systematic empirical investigation of scaling behaviors in RL-based post-training, with a particular focus on mathematical reasoning. Based on a set of experiments across the full Qwen2.5 dense model series (0.5B to 72B), we characterize how model scale, data volume, and computational budget interact to shape performance. Our analysis leads to four key findings: ❶ Larger models consistently exhibit superior learning efficiency on both compute and data metrics. ❷ The relationship between test loss, compute, and data can be modeled by a predictive power-law which is robust across both base and instruction-tuned models. ❸ Although larger models exhibit higher learning efficiency, the analytical learning efficiency term $k(N)$ in the power-law reveals a latent saturation trend in learning efficiency as model size continues to increase. ❹ In data-constrained regimes, repeated reuse of high-quality data proves highly effective, as final performance is primarily governed by the total number of optimization steps rather than the uniqueness of samples. Collectively, these results provide a principled foundation and practical guidelines for efficiently scaling the reasoning capabilities of LLMs through RL post-training.
1 Introduction
The rapid progress of large language models (LLMs) has made elucidating their scaling laws a matter of central importance. These laws, which capture the intricate relationships between model architecture, parameter size, computational cost, data availability, and downstream performance (kaplan2020scaling; hoffmann2022training), are invaluable not only because they illuminate the latent factors governing learning dynamics, but also because they provide actionable guidance on how to distribute scarce computational resources most effectively (li2025misfitting). While extensive efforts have clarified scaling behavior, the scaling behavior of reinforcement learning (RL) post-training for LLM reasoning remains underexplored.
During pretraining, kaplan2020scaling show that cross-entropy loss follows smooth power-law scaling in model size, dataset size, and training compute, implying that larger models trained for fewer steps are compute-optimal. hoffmann2022training refine this by showing that, under fixed compute, scaling parameters and tokens proportionally is optimal, since many large models are undertrained.Extending to neural-based RL, hilton2023scaling empirically demonstrates that the intrinsic performance of convolutional neural networks (CNNs) optimized via reinforcement learning also scales like power-law with model capacity and environment interaction.
Recently, RL has become the predominant post-training strategy for enhancing the reasoning abilities of LLMs, particularly in mathematics, a domain that demands long-horizon, compositional reasoning (ferrag2025llm; guo2025deepseekr1; kimi2025scaling; ahn2024large). Given the recent works applying RL for LLM reasoning, it is crucial to understand how to scale RL training effectively for reasoning tasks. In this work, we present a systematic empirical study to characterize these scaling behaviors. We focus on three critical resource regimes: (1) the compute-constrained scenario, where we seek to minimize test loss under a fixed FLOPs budget by determining the optimal model size; (2) the data-constrained scenario, where we identify the model size that yields the lowest test loss given a limited number of unique samples; and (3) the data reuse scenario, where we investigate the trade-off between unique data volume and repeated optimization steps under a fixed computation budget.
Based on our analysis across these regimes, we propose a predictive formulation that characterizes the relationship between test loss $L$, model size $N$, and resource budget $X$ (where $X$ denotes either Compute $C$ or Data $D$). We find that the scaling behavior can be effectively modeled by a power law that exhibits a log-linear relationship between test loss and resource budget:
Core Scaling Formula $\log L(N,X)=-k(N)\cdot\log X+E(N).$ (1)
Crucially, our empirical observations suggest that the learning efficiency $k(N)$ does not grow indefinitely with the increasing model scale. Instead, it follows a saturation trend and can be modeled by the formula(Eq. 2) below:
RL Post-Training Learning Efficiency $k(N)=\frac{K_{\text{max}}}{1+\frac{N_{0}}{N}}.$ (2)
This formulation shows that: while larger models consistently exhibit higher learning efficiency (represented by $k(N)$), the marginal gains in efficiency diminish as model size increases, asymptotically approaching a theoretical limit $K_{\text{max}}$.
To empirically validate this law and determine its parameters, we fine-tune 63 LLMs with reinforcement learning on 50k mathematics problems, based on the Qwen2.5 model family (qwen2025qwen25technicalreport). Figure 1 shows that, within the 0.5B-72B range, the loss reduction brought by RL follows an approximately log-linear trend with compute and data size. Importantly, larger models not only have better initial performance but also generally have more efficiency in computation and data utilization during the optimization process. Through deeper analysis, we find that the scaling behavior exhibits notable predictability and further reveal that efficiency gains from model scaling diminish gradually, eventually leading to a saturation effect for larger models.
We further analyze the data-constrained regime, where we demonstrate that data reuse is a highly effective strategy. We validate the generality of our findings through extensive ablation studies on both base and instruct model series. Besides, we have also conducted experiments to study the impact of the rollout number in the GRPO algorithm (shao2024deepseekmathpushinglimitsmathematical), shown in Appendix B.2. These investigations establish fundamental scaling relationships for RL post-training, providing a quantitative foundation and practical guidelines for resource-efficient model refinement.
(a) Relation between test loss and post training compute for base models, fitted on 0.5B-32B and extrapolated on 72B.
(b) Relation between test loss and post training data size for base model, fitted on 0.5B-32B and extrapolated on 72B.
Figure 1: Training Data Fiited on 0.5B-32B with Extrapolation on 72B: In both cases, larger models consistently exhibit higher learning efficiency than smaller models.
Specifically, our key findings can be summarized as follows:
  • In our experiment scale, larger models, starting with stronger initial performance, consistently achieve better compute and data efficiency in RL post-training for mathematical reasoning. However, the marginal gains in this efficiency diminish gradually, revealing a saturation trend as model scale increases.
  • The scaling law exhibits predictive capability across both base and instruction-tuned models, allowing us to forecast the training efficiency of larger models and predict the remaining training trajectory from early training data.
  • In data-limited settings, repeated exposure to a small dataset is nearly as effective as using larger corpora, highlighting data reuse as a practical strategy.
2 Experimental Setup
We describe the experimental setup for studying scaling behavior in RL post-training of LLMs for mathematical reasoning, including the model family, training and evaluation data, and evaluation protocol in this section. Full details are provided in Appendix A.
Models and Framework.
We use the Qwen2.5 model family (0.5B, 1.5B, 3B, 7B, 14B, 32B and 72B parameters) (qwen2025qwen25technicalreport), which shares the same architecture, so that parameter count is the only variable in our scaling analysis. All experiments are run with the VeRL framework (sheng2024hybridflow), a large-scale RL platform for LLMs ensuring consistency and reproducibility.
Dataset settings.
The training data is the mathematics subset of the guru-RL-92k dataset from the Reasoning360 project (cheng2025revisiting), which is carefully curated through deduplication and difficulty filtering. We further sort the problems by increasing difficulty (decreasing pass rate, evaluated by Qwen2.5-7B-Instruct model) to enable curriculum learning. The evaluation data consists of two parts. To derive scaling laws, we use a held-out set of 500 in-domain math problems sampled from the training distribution. To assess generalization, we evaluate on a broader benchmark suite spanning mathematics (AIME2024 (patel2024aimeaioptimizationmultiple), AMC2023 (knoveleng_amc23_dataset), GSM8K (cobbe2021trainingverifierssolvemath), MATH500 (lightman2023letsverifystepstep)), code (HumanEval (chen2021evaluatinglargelanguagemodels)), logic (Zebra Puzzle(Lin_ZeroEval_A_Unified_2024)), and science (SuperGPQA (pteam2025supergpqascalingllmevaluation)). More details about dataset settings can be found in Appendix A.1.
Prompt Setting.
To ensure stable behavior during RL training and evaluation, we use structured prompts tailored to each domain. For example, all mathematics problems are prepended with the Chain-of-Thought prompt (wei2023chainofthoughtpromptingelicitsreasoning): “You are a knowledgeable math assistant. Answer the following questions and think step by step”. More prompt templates for all related domains could be found in Appendix A.3.
RL Algorithm.
[…]
$\mathcal{L}_{\text{GRPO}}=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{\lvert o_{i}\rvert}\sum_{t=1}^{\lvert o_{i}\rvert}\Bigl\{\min\Bigl[\rho(\theta)\,\hat{A}_{i,t},\;\operatorname{clip}\!\Bigl(\rho(\theta),1-\varepsilon,\,1+\varepsilon\Bigr)\hat{A}_{i,t}\Bigr]-\beta\,\mathrm{D}_{\mathrm{KL}}\Bigr\},$ (3)
where $\rho(\theta)=\frac{\pi_{\theta}\!\left(o_{i,t}\mid q,\,o_{i,<t}\right)}{\pi_{\theta_{\mathrm{old}}}\!\left(o_{i,t}\mid q,\,o_{i,<t}\right)}$ is the important sampling weight. For each output $o_{i}$, a reward model or rule is used to yield the reward signal $\mathbf{r}=\{r_{1},r_{2},\cdots,r_{G}\}$. The advantage is computed as
$\hat{A}_{i,t}=\frac{r_{i}-\mathrm{mean}(\mathbf{r})}{\mathrm{std}(\mathbf{r})}.$ (4)
Evaluation Process.
[…]
Metric.
Our primary evaluation metric is the test loss ($L$), a proxy for reward-based performance in the RL setting. Formally, $L=1-(R/R_{\text{max}})$, where $R$ is the number of correct solutions and $R_{\text{max}}$ the total. We adopt the term "test loss" for consistency with foundational neural scaling law literature (kaplan2020scaling). Notably, maximizing reward in RL training is equivalent to minimizing $L$.
Fitting and Prediction Protocols.
To systematically evaluate the robustness and predictive capability of our derived scaling laws, we employ two distinct fitting protocols throughout our analysis:
  • Inter-model Extrapolation: We fit the scaling law parameters using data from smaller models (0.5B to 32B) to calculate the learning efficiency and predict the performance of the larger model (72B).
  • Intra-model Extrapolation: We fit the scaling law using only the early training steps of a specific model to forecast its loss trajectory for the remainder of the training process.
3 Empirical Results and Scaling Laws
This section presents a comprehensive empirical investigation into the scaling behavior of RL for post-training LLMs. We first examine scaling behaviors under compute and data constraints, then analyze independent scaling dimensions, data reuse strategies, and finally evaluate generalization performance together. To ensure robust conclusions, each configuration is repeated three times for both base and instruct models ranging from 0.5B to 72B.Their statistical uncertainty analysis, including Average Standard Deviation and Standard Error of the Mean (SEM), are provided in Appendix C.3.
3.1 Compute-Optimal Scaling
To characterize the scaling behavior under computational limits, we first formalize the Compute-Constrained Scenario. Given a fixed computational budget $C$, we seek to identify the optimal model size $N$ (and the corresponding data allocation $D$) that minimizes the final test loss. This can be expressed as the following constrained optimization problem:
$\underset{N,~D}{\operatorname*{arg\,min}}\;L(N,D)\quad\text{s.t.}\quad\mathrm{FLOPs}(N,D)=C_{\text{const}},$ (5)
Observation 1 Within the 0.5B–32B parameter range, RL post-training achieves lower test loss under a fixed budget $C$ by prioritizing larger models over extended training duration. However, between 32B and 72B, learning efficiency saturation reduces scaling’s marginal gains, introducing a trade-off between model scale and training steps.
(a) Test loss vs training compute with extrapolation on 72B for base model
(b) Test loss vs training compute with extrapolation on 72B for instruct model
(a) Test loss vs training compute with extrapolation for remainder of each training process on base model
(b) Test loss vs training compute with extrapolation for remainder of each training process on instruct model
Figure 3: Compute Scaling and Predictive Capability from 0.5B-72B for Base and Instruct Models
In the compute-constrained setting, we train 0.5B–72B models and measure test loss as a function of cumulative FLOPs $C$. As shown in Figure 3, larger models consistently outperform smaller ones under the same compute budget for both base and instruct variants. These plots include both Inter-model Extrapolation (fitted on 0.5B-32B and extrapolated on 72B) and Intra-model Prediction (predicting the remainder of training from initial steps) to demonstrate the predictive power of our derived scaling law. The loss–compute relationship follows a log-linear trend, which can be modeled by a power law:
$\log(L(N,C))=-k_{C}(N)\cdot\log(C)+E_{C}(N),\ \text{where}\ \ k_{C}(N)=\left(\frac{K_{Cmax}}{1+\frac{N_{C}}{N}}\right)$ (6)
To demonstrate the predictive capability of the proposed formula Eq 6, we evaluate it in two distinct extrapolation settings:
  1. 1.
    Inter-model Extrapolation: We fit the law’s parameters on smaller models (0.5B–32B) to calculate the learning efficiency ($k_{C}(N)$) of 72B model. As shown in Figure 2(a) and 2(b), the predicted efficiency aligns closely with the actual 72B performance.
  2. 2.
    Intra-model Prediction: We fit the law using only early training steps to forecast the remaining trajectory for a specific model, shown in Figure 3(a) and 3(b).
We further analyze learning efficiency term $k_{C}(N)$ in Eq. 6. As Figure 4 shows, $k_{C}(N)$ grows with model size $N$, meaning larger models consistently have higher learning efficiency. However, the efficiency gain from model scale is not uniformly linear. Beyond 32B, the increase in $k_{C}(N)$ diminishes, leading to efficiency saturation.
Figure 4: Fitted learning efficiency $k_{C}(N)$ for Base(Left) and Instruct(Right) models: Larger models consistently exhibit higher learning efficiency with efficiency gains begin to diminish after 32B model.
This saturation is manifested as a distinct performance crossover in Figure 3: In contrast to the immediate dominance of larger models in smaller parameter regimes, the 32B model outperforms the 72B counterpart initially under equivalent compute budgets, as the smaller model size inherently enables more training steps. We believe this observation reveals a latent trade-off between model scale and training steps in compute-constrained scenarios.
3.2 Data-Optimal Scaling
In many practical applications, the bottleneck lies not in compute but in the availability of high-quality reasoning data. We define the Data-Constrained Scenario as determining the model size $N$ that yields the lowest test loss given a limited amount of unique training data $D$:
$\underset{N,~C}{\operatorname*{arg\,min}}\;L(N,C)\quad\text{s.t.}\quad D=D_{\text{const}},$ (7)
Observation 2 Within the model range of 0.5B to 72B, for a fixed volume of unique training data $D$, larger models demonstrate superior sample efficiency, consistently achieving lower test loss.
(a) Test loss vs data size with extrapolation on 72B for base model
(b) Test loss vs data size with extrapolation on 72B for instruct model
(a) Test loss vs data size with extrapolation for remainder of each training process on base model
(b) Test loss vs data size with extrapolation for remainder of each training process on instruct model
Figure 6: Data Scaling and Predictive Capability from 0.5B-72B for Base and Instruct Models
In the data-constrained setting, we train models with varying parameter counts $N$ on fixed amounts of unique samples $D$. As shown in Figure 6, larger models consistently achieve lower test loss and higher sample efficiency across both base and instruct variants. The loss–data relationship also follows a log-linear trend and can be modeled by a similar formula addressed in compute scenario:
$\log(L(N,D))=-k_{D}(N)\cdot\log(D)+E_{D}(N),\ \text{where}\ \ k_{D}(N)=\left(\frac{K_{Dmax}}{1+\frac{N_{D}}{N}}\right)$ (8)
Mirroring the analysis in Section 3.1, we evaluate the extrapolative capability of our data scaling law (Eq. 8) in two settings:
  1. 1.
    Inter-model Extrapolation: By fitting parameters on smaller models (0.5B–32B), we accurately predict the data efficiency ($k_{D}(N)$) on 72B model, as illustrated in Figure 5(a) and 5(b).
  2. 2.
    Intra-model Prediction: We forecast the loss trajectory for the remainder of the training process using only early-stage data, shown in Figure 6(a) and 6(b).
We adopt the same analytic form for the data efficiency coefficient $k_{D}(N)$ as we did for compute. As illustrated in Figure 7, $k_{D}(N)$ follows a saturation curve identical to $k_{C}(N)$: while larger models excel at extracting knowledge from each data point, the efficiency gains diminish at scales beyond 32B. The unified functional form across both compute and data domains reflects the theoretical consistency of our scaling law.
Figure 7: Fitted learning efficiency $k_{D}(N)$ for Base(Left) and Instruct(Right) models: $k_{D}(N)$ exhibit nearly identical growth trends with $k_{C}(N)$, with efficiency gains also starting to diminish beyond the 32B model.
3.3 Scaling up Model Size
Observation 3 When trained to converge on sufficiently large datasets, test loss decreases monotonically with model size, though the trend deviates from a strict power law.
(a) Fig (a) shows Relation between test loss and model size NN for our trained model and SOTA models demonstrates the effectiveness of our training.
(b) Fig (b) shows relation between response length and data size for instruct models,larger models generate longer responses except for 32B model.
Figure 8: Analysis of model scaling properties shows the effectiveness of our training process in this empirical study.
We train models of varying sizes to convergence and compare their final test loss. As shown in Figure 8(a), larger models consistently achieve lower loss, improving monotonically with scale. The curve deviates from a strict power law: smaller models show weaker gains, suggesting diminishing returns at low parameter counts. A likely reason is that larger models inherit richer pre-trained representations, which reinforcement fine-tuning exploits for greater improvements than parameter growth alone would predict. Figure 8(b) further shows that as RL training progresses, larger models generate longer responses except for 32B model. This correlates with higher accuracy, indicating greater test-time scaling efficiency: additional inference tokens yield larger gains in bigger models.
We also benchmark our RL-tuned Qwen2.5 models (qwen2025qwen25technicalreport) against state-of-the-art open-source reasoning systems, including Qwen3 (yang2025qwen3technicalreport) and GPT-OSS (openai2025gptoss120bgptoss20bmodel), detailed in Table 4. On our held-out set, the 32B and 72B models match or surpass dense Qwen3 counterparts of similar size, highlighting the effectiveness of RL post-training. Mixture-of-experts models such as Qwen3 and GPT-OSS achieve approximate loss at much larger scales (235B), with GPT-OSS-120B currently leading. These comparisons suggest that scaling across 0.5B-72B will be necessary to fully characterize post-training behavior and compete with frontier MoE systems.
3.4 Scaling with Constrained Data and Reuse
When unique data is scarce, a critical question is whether repeating data is effective. We investigate this Data Reuse Scenario by fixing the total data budget and varying the reuse factor $\tau$. Specifically, we aim to identify the optimal reuse factor $\tau$ that minimizes the test loss:
$\underset{\tau}{\text{argmin}};L(\tau)\quad\text{s.t.}\quad D_{\text{unique}}\times\tau=D_{\text{total}},$ (9)
Observation 4 Performance in data-constrained settings is primarily determined by the total number of used data ($D_{\text{total}}$). For a fixed $D_{\text{total}}$, the final test loss is remarkably insensitive to the data reuse factor ($\tau$), with no significant degradation up to $\tau=25$.
(a) Base Model
(b) Instruct Model
Figure 9: Plot shows the final test loss of a 7B Base (left) and Instruct (right) model trained with a fixed toal number of samples but varying the data reuse factor $\tau$.
We further consider the data-reuse scenario, where high-quality data is limited but can be revisited multiple times. To simulate this, we partition the training set into smaller subsets while preserving the difficulty distribution (Details provided in Appendix A.4). Each subset is cycled through multiple times, with the reuse factor $\tau$ controlling how often each unique example is revisited. The total number of used data $D_{\text{total}}$ is fixed across runs, and curriculum ordering is maintained so that problems are always presented from easy to hard. This ensures that performance differences arise solely from the degree of data reuse, rather than distributional or scheduling artifacts.
As shown in Figure 9, performance remains nearly unchanged for $\tau\leq 25$, while moderate degradation appears as $\tau$ increases further. At $\tau=100$, we observe clear signs of overfitting, indicating that repeated reuse eventually harms generalization. Overall, these results suggest that final performance is primarily governed by the total number of used data rather than sample uniqueness, and that moderate data reuse is an effective strategy for RL fine-tuning with limited datasets.
3.5 Domain Transfer
Observation 5 RL post-training on mathematical reasoning yields generalization improvements on in-domain tasks with varying difficulty, but shows negligible transfer to out-of-domain tasks.
Figure 10: The effect of domain transfer, illustrated with the Qwen2.5-72B-Instruct model.
We also investigate the generalization capabilities of the reinforcement learning fine-tuning (RFT) models by evaluating them on a suite of unseen in-domain tasks with varying difficulty and out-of-domain (OOD) tasks. More results are in Appendix B.1.
In-Domain Generalization. Figure 10 shows consistent improvements on unseen mathematics tasks outside the training set. On benchmarks, from easy to hard, including GSM8K, MATH-500, AMC2023, AIME2024, test loss steadily decreases with training compute, suggesting that RL post-training enhances transferable reasoning skills within mathematics.
Out-of-Domain Generalization. As shown in Figure 10, results on OOD tasks are markedly different. For code generation (HumanEval) and STEM problems (SuperGPQA), performance gains marginally, indicating that RL fine-tuning is highly specialized.On logical reasoning (zebra_puzzle), performance degrades for larger models, suggesting that intensive optimization on mathematical reasoning may interfere with or "damage" other distinct reasoning abilities.
4 related work
Foundational Scaling Laws of Neural Language Models. Foundational scaling studies show language modeling loss follows smooth power-laws in model size $N$, data $D$, and compute $C$ (kaplan2020scaling), with compute-optimal training prescribing near lockstep growth of parameters and tokens under fixed FLOPs (hoffmann2022training). Later analyses attribute earlier discrepancies to embedding/non-embedding parameter accounting, last-layer costs, optimizer warmup, and scale-sensitive hyperparameters (pearce2024reconciling; porian2024resolving), while data-centric refinements examine pruning efficiency (sorscher2022beyond), repetition effects (hernandez2022scaling), gzip-based complexity predictors (pandey2024gzip), constrained or synthetic regimes (muennighoff2023scaling; qin2025scaling), and task transfer (e.g., translation) (isik2024scaling). Test-time compute amplification supplies an inference analogue to classical training laws (snell2024scaling).
RL post-training in LLMs. In RL, power-law trends similarly link capacity, interaction compute, and performance (hilton2023scaling); scaling RFT across horizon and compute improves mathematical and coding reasoning (guo2025deepseekr1; kimi2025scaling; mai2025agent; zhang2025landscapeagenticreinforcementlearning; zhang2025surveyreinforcementlearninglarge), while extended schedules (liu2025prorl), ultra-low-shot or single-example RL (wang2025reinforcement), and minimal-data efficiency paradigms (li2025limr) probe data–compute tradeoffs. Instability and uneven gains highlight fragile optimization (zeng2025simplerlzoo; yue2025doesrlincentivize), and multi-domain mixtures reveal both synergy and interference across math, code, and logic (li2025onedomainhelps; cheng2025revisiting).
Mathematical Reasoning with LLMs. Mathematical reasoning amplifies these dynamics: accuracy generally scales upward while verification behaviors remain inconsistent (touvron2023scaling); corpus volume and quality jointly shape attainable curves (ye2024skywork); multi-task math-generalist training diverges from specialist scaling trajectories (yue2023mammoth); and RL with code execution induces additional behaviors such as emergent tool use concentrated in math problem solving (zeng2025agent). Collectively, evidence indicates that reasoning performance is governed by interacting axes of model size, data distribution/quality, training (supervised vs. RL) paradigm, and allocation of both training and inference compute, while unified laws for mathematical reasoning remain only partially characterized.
5 Discussion
Scaling Dependence on Evaluation Environment and Metrics. Reinforcement learning optimizes directly for environment rewards (10.5555/3312046), which in principle allows unbounded capability—as demonstrated by AlphaZero mastering board games (silver2017masteringchessshogiselfplay), AlphaFold predicting protein structures (jumper2021highly), and frontier LLMs such as Gemini-2.5-Pro achieving IMO-level performance (huang2025gemini25procapable). In contrast, text-based LLMs lack a well-defined RL environment, forcing us to rely on human-curated datasets as proxies. Test loss thus serves as a pragmatic but imperfect metric: it is monotonic and convergent, yet heavily dependent on dataset construction and task difficulty, with different benchmarks (e.g., GSM8K vs. AIME, Section 3.5) showing distinct convergence rates. This task dependence makes the absolute coefficients of our fitted scaling laws ($k(N),E$) difficult to interpret universally. Prior work proposed “intrinsic performance”—the minimum compute needed to reach a target reward—as a normalization across environments (hilton2023scaling), but we did not find an analogous measure in large-scale LLMs. Establishing principled, environment-independent evaluation protocols remains an open and critical challenge for RL-based scaling studies.
Scaling Dependence on Model Scale. Our study of models from 0.5B to 72B parameters shows that larger models exhibit greater sample and compute efficiency in RL post-training. This parameter range allows us to characterize the scaling limits. We found that these advantages do not extend indefinitely. Our analytic learning efficiency term $k(N)$ in Eq.6 and Eq.8, explicitly confirms that the efficiency gains follow a saturation curve toward a limit ($K_{\max}$). This finding implies that scaling up models beyond a certain point, while still yielding absolute performance gains, suffers from diminishing marginal returns in efficiency.
Dependence on RL Algorithm. Our analysis is based on GRPO, a mainstream and stable RL post-training algorithm for LLMs that uses an actor-only design and normalizes rewards across responses. Comparative study with alternative RL algorithms  (cui2025entropymechanismreinforcementlearning) reports minor differences in training curves. Whether more advanced algorithms can significantly improve sample efficiency or stability—and thereby reshape the scaling frontier—remains an important open question.
Future of LLM Agent. The integration of reinforcement learning with agentic LLMs is increasingly viewed as a promising direction (zhang2025landscapeagenticreinforcementlearning; zhang2025surveyreinforcementlearninglarge). Both theoretical and empirical studies show that augmentations such as external tool use and long-term memory can substantially boost model performance (lin2025understandingtoolintegratedreasoning; houliston2025provablebenefitsintoollearning; mai2025agentrlscalinglaw). We anticipate that such agentic mechanisms will markedly improve the scaling behavior of RL-trained LLMs: by offloading deterministic computations to tools and focusing learning on high-level decision making, these models could achieve much higher efficiency, effectively shifting the performance frontier upward for a given compute or data budget. Understanding the scaling laws of these agentic systems is, therefore, a key and exciting avenue for future research.
6 Conclusion
This study presents the systematic exploration of scaling laws for reinforcement learning post-training of large language models in mathematical reasoning. Through 54 controlled experiments across the Qwen2.5 family, we show that larger models consistently achieve superior compute and data efficiency, that performance in data-limited regimes depends primarily on the total number of optimization steps rather than data uniqueness, and that moderate data reuse can be highly effective without harming generalization. While RL post-training reliably strengthens in-domain reasoning, its transfer to out-of-domain tasks remains limited, underscoring the trade-off between specialization and breadth. Our ablation further identifies rollout size in GRPO as a practical lever tied to compute budgets. Taken together, these findings offer principled and actionable guidelines for resource-efficient RL fine-tuning and suggest promising directions for further exploration of scaling and generalization in LLM reasoning.
Appendix A Experiment Setup Details
This section provides a detailed breakdown of the datasets and hyperparameters used in our experiments, supplementing the information provided in the main text.
A.1 Dataset Details
Our training was conducted on a curated mathematics dataset. For evaluation, especially for analyzing generalization (as mentioned in the main text), we utilized a comprehensive suite of benchmarks spanning multiple domains. The composition of this evaluation suite is detailed in Table 1.
Table 1: Composition of the multi-domain evaluation suite.
Dataset Samples Huggingface Tag Domain
Held-out Data 500 LLM360/guru-RL-92k Math
aime2024 30 Maxwell-Jia/AIME_2024 Math
amc2023 40 knoveleng/AMC-23 Math
codegen_humaneval 164 openai/openai_humaneval Code
gsm8k 1319 openai/gsm8k Math
logic_zebra_puzzle 200 LLM360/guru-RL-92k Logical Reasoning
math 500 HuggingFaceH4/MATH-500 Math
stem_supergpqa 200 LLM360/guru-RL-92k STEM
Total 2953
A.2 Hyperparameter Configuration
All experiments were conducted with a consistent set of hyperparameters for the Group Relative Policy Optimization (GRPO) algorithm to ensure a fair comparison across different model sizes and configurations. The key hyperparameters are listed in Table 2.
Table 2: GRPO training hyperparameters used across all experiments.
Hyperparameter Value
Learning Rate $1.0\times 10^{-6}$
Batch Size 512
KL Loss Coefficient 0.001
Rollout Temperature (Training) 1.0
Rollout Temperature (Evaluation) 0.7
Clip Ratio (High & Low) 0.2
Input Sequence Length 2048
Output Sequence Length 4096
A.3 Prompt Templates
This section details the specific prompt templates used for evaluating models on different domains. For each task, the model was provided with the corresponding instruction prepended to the problem statement <question>.
Table 3: Prompt templates used for different evaluation domains.
Domain Prompt Template
[…]
A.4 Data Reuse Experiment Setup
Figure 11: Data Reuse Schema
To systematically evaluate the effect of data reuse under constrained data scenarios, we design controlled experiments where all runs are trained with the same total data size but different levels of data repetition. Each run randomly samples a subset from the full training corpus and repeats this subset sufficiently many times to exactly match the target data budget (i.e., $\text{subset size}\times\tau=\text{total data size}$). Unlike muennighoff2023scaling, subsets are sampled independently for each run rather than sampling within the larger subsets, to mitigate sampling bias and balance stochasticity across conditions. To remain consistent with the Curriculum Learning setting of the main experiments, examples within each subset are ordered by increasing difficulty; across epochs, this difficulty schedule is preserved and repeated rather than reshuffled, as illustrated in Figure 11.
Appendix B Additional Experiment Results
This section provides supplementary experimental results that support and extend the analyses presented in the main body of the paper.
B.1 Performance on In-Domain and Out-of-Domain Tasks
To assess how the mathematical reasoning capabilities acquired during RL fine-tuning generalize, we evaluated our models on a comprehensive suite of unseen benchmarks. We categorize these into two groups: in-domain different tasks (other mathematics datasets) and out-of-domain tasks (e.g., code, science, logic). The results are presented in Figure 12 and Figure 13.
Figure 12: Test loss on in-domain and out-of-domain benchmarks vs data size for Base models. It shows modest positive transfer on in-domain tasks, with limited or negative transfer on OOD tasks.
In-Domain Generalization (Different Mathematical Tasks). On mathematics benchmarks not included in our training set (such as GSM8K, MATH, AIME, and AMC), we observe a generally positive transfer of learned skills. For most of these tasks, the test loss shows a modest but consistent decrease as training progresses, particularly for the larger models. This suggests that the model’s enhanced reasoning ability is not overfitted to the training distribution and is applicable to a wider range of mathematical problems.
Out-of-Domain Generalization. When evaluating on tasks outside of mathematics, the generalization is more limited. For both code generation (HumanEval) and science problems (SuperGPQA), performance remains largely static throughout the training process across all model sizes, with test loss curves staying flat. This indicates that the specialized mathematical reasoning skills do not readily transfer to these domains. A noteworthy phenomenon is observed in the logical reasoning task (Zebra Puzzle): the largest models (particularly the 14B variants) show a degradation in performance (an increase in test loss) as training progresses, suggesting a potential negative transfer effect where intensive optimization on mathematical reasoning may interfere with capabilities required for certain types of logical puzzles.
Figure 13: Test loss on in-domain and out-of-domain benchmarks vs data size for Instruct models.
B.2 Ablation on GRPO Hyperparameters
(a) 7B-Base: Loss vs. Compute
(b) 7B-Base: Loss vs. Data Size
(c) 7B-Instruct: Loss vs. Compute
(d) 7B-Instruct: Loss vs. Data Size
Figure 14: Effects of GRPO rollout size on training efficiency
We conducted an ablation study on the rollout group size $G$, a key GRPO hyperparameter that controls how many responses are sampled per prompt. This directly affects both the compute per update and the stability of the training signal. We tested $G\in\{4,8,16,32\}$ on the 7B models.
Data-centric View. Figure 14(b) and 14(d) shows that larger rollout sizes consistently yield better sample efficiency: $G=32$ achieves the lowest test loss for the same number of unique samples. This supports the intuition that more responses per question provide a stronger advantage estimate and thus more effective gradient updates.
Compute-centric View. The optimal rollout size $G$ is not fixed but shifts with the training budget. This implies that practitioners should tune $G$ according to available compute rather than relying on a universal setting. We attribute this dynamic to the trade-off between the higher variance reduction from larger $G$ and the additional FLOPs it consumes, which makes small $G$ preferable at low budgets but large $G$ superior when ample compute is available.
B.3 Peformance Compared with Advance Models
Model Family Model Identifier Pass@1 Score
Models from Our Study (Post-RFT)
Qwen2.5-Base 0.5B 0.070
1.5B 0.116
3B 0.182
7B 0.338
14B 0.450
32B 0.540
72B 0.607
Qwen2.5-Instruct 0.5B 0.078
1.5B 0.138
3B 0.216
7B 0.380
14B 0.488
32B 0.590
72B 0.617
External SOTA Models (for Comparison)
Qwen3 0.6B 0.178
1.7B 0.288
4B 0.418
8B 0.366
14B 0.388
30B (A3B) 0.528
32B 0.412
235B (A22B) 0.602
GPT-OSS 20B 0.556
120B 0.660
Table 4: Performance of various models on the held-out evaluation set.
To contextualize the performance of our models and the difficulty of our primary evaluation metric, we benchmarked a range of external, state-of-the-art (SOTA) models on our held-out mathematics test set. The results are presented in Table 4. The performance of our Qwen2.5 models reflects their final scores after the completion of reinforcement learning fine-tuning (RFT), while others are benchmarked directly.
Appendix C Formula Fitting and Derivation
C.1 FLOPs Calculation Methodology
The computational cost for a LLM is primarily determined by the number of non-embedding parameters ($N$) and the number of processed tokens ($T$). The costs for the fundamental operations are:
  • Forward Pass Cost: The cost of a single forward pass is approximately $C_{\text{fwd}}\approx 2NT$ FLOPs.
  • Backward Pass Cost: The backward pass is approximately twice as expensive as the forward pass, so $C_{\text{bwd}}\approx 4NT$ FLOPs.
A full training step, which includes one forward and one backward pass for the gradient update, therefore has a total computational cost of:
$C_{\text{train}}=C_{\text{fwd}}+C_{\text{bwd}}\approx 2NT+4NT=6NT\text{ FLOPs.}$ (10)
$\text{FLOPs}_{\text{step}}=6\times N\times T_{\text{step}}$ (11)
By recording the exact number of processed tokens $T$ per step, we compute the cumulative FLOPs reported throughout this paper as the sum of these per-step calculations over the course of training.
C.2 Coefficient Comparison
We consider the two laws
$\ln L(N,C)=-\,k_{C}(N)\,\ln C+E_{C}(N),$ (12)
and
$\ln L(N,D)=-\,k_{D}(N)\,\ln D+E_{D}(N),$ (13)
are consistent under the linkage $C=ND\phi$ where $\phi>0$ is a constant for simplification.
Claim. Under $C=ND\phi$, the slopes coincide and the intercepts differ by a known shift:
$k_{C}(N)$ $=k_{D}(N)\;=\;k(N),$ (14)
$E_{C}(N)$ $=E_{D}(N)\;+\;k(N)\,\ln\!\bigl(N\phi\bigr).$ (15)
Proof. Substitute $C=ND\phi$ into equation 12:
$\ln L(N,C)$ $=-\,k_{C}(N)\,\ln\!\bigl(ND\phi\bigr)+E_{C}(N)$
$=-\,k_{C}(N)\,\bigl[\ln D+\ln(N\phi)\bigr]+E_{C}(N)$
$=-\,k_{C}(N)\,\ln D\;+\;\Bigl(E_{C}(N)-k_{C}(N)\,\ln(N\phi)\Bigr).$
Comparing this with equation 13, i.e., $\ln L(N,D)=-\,k_{D}(N)\,\ln D+E_{D}(N)$, equality for all $D>0$ forces the coefficients of $\ln D$ and the constants to match:
$k_{D}(N)=k_{C}(N)=:k(N),\qquad E_{D}(N)=E_{C}(N)-k(N)\,\ln(N\phi).$
Rearranging the second identity yields equation 15. $\square$
The observation from Figure 4 and Figure 7 also matches with this conclusion.
C.3 Fitting for K and E
Table 5: Uncertainty Analysis for raw data: Base and Instruct Models (Holdout Score)
Model Base Instruct
Test Loss Avg Std SEM Test Loss Avg Std SEM
0.5B 0.9419 0.0082 0.0048 0.9458 0.0073 0.0042
1B 0.9129 0.0091 0.0053 0.8988 0.0098 0.0057
3B 0.8582 0.0129 0.0074 0.8281 0.0112 0.0065
7B 0.7148 0.0147 0.0085 0.6777 0.0142 0.0082
14B 0.6051 0.0149 0.0086 0.5588 0.0143 0.0083
32B 0.4937 0.0056 0.0032 0.4579 0.0127 0.0073
72B 0.4359 0.0143 0.0082 0.4320 0.0140 0.0081
Table 6: Comparison of $k_{max}$ and $N_{0}$ Parameters Across Fitting Scenarios
Source Metric Scenario $k_{max}$ $N_{0}$ (B) $R^{2}$
Base L(N,C) Intra-model 0.1349 13.09 0.9955
L(N,C) Inter-model 0.1518 17.37 0.9944
L(N,D) Intra-model 0.1348 11.52 0.9953
L(N,D) Inter-model 0.1631 16.95 0.9947
Instruct L(N,C) Intra-model 0.1276 17.27 0.9970
L(N,C) Inter-model 0.1443 28.33 0.9950
L(N,D) Intra-model 0.1325 17.08 0.9970
L(N,D) Inter-model 0.1484 27.15 0.9949
Appendix D A Loss Decomposition Model for Scaling Analysis
During the analysis, we found a more generalized form of the potential scaling law function that fits the curves well. This model fits the same dataset as the main experiments and is included here as a formally documented alternative for future research.
D.1 Loss Decomposition Model
The General Loss Decomposition.
We construct the generalized formula as follows, based on the observation of the loss composition of post-training and the experiment data:
$L(N,D)=L_{\infty}+G(N)+\lambda(N)\cdot P(N,D)$ (16)
Each term in Equation equation 16 represents a clear part decomposing the loss:
  • $L_{\infty}$ denotes the irreducible loss, representing the fundamental loss floor that persists even with infinite model capacity and unlimited data. It reflects task-intrinsic uncertainty and noise that cannot be eliminated by improved modeling or additional training, such as inherent stochasticity in the environment or irreducible mismatch between training and evaluation distributions.
  • $G(N)$ denotes the model-limited loss, capturing the asymptotic loss floor imposed by finite model capacity $N$ in the limit of infinite data. It corresponds to the capacity-dependent performance frontier of models with size $N$.
  • $\lambda(N)$ denotes the learnable capacity, defined as the maximum achievable reduction in loss that a model of size $N$ can attain through post-training, beyond its model-limited loss. In RL post-training settings, this term can depend on the pretraining regime, as well as the degree of mismatch between the pretraining and post-training task settings. Conceptually, the learnable capacity is governed by two opposing effects: larger models generally have greater capacity to extract and acquire new knowledge, while simultaneously having already absorbed more information during pretraining, leaving less headroom for additional improvement via RL. As a result, the monotonic dependence of $\lambda(N)$ on model size is non-trivial, and its precise modeling likely requires additional assumptions and empirical characterization.
  • $P(N,D)$ denotes the learning progress, a normalized function taking values in $[0,1]$ that quantifies the fraction of the learnable capacity $\lambda(N)$ realized when training on a dataset of size $D$.
Instantiated Model.
Following prior empirical scaling law studies, we parameterize the model-limited loss as
$G(N)=\left(\frac{N_{0}}{N}\right)^{\alpha},$ (17)
reflecting the observed power-law dependence of loss on model size in the infinite-data regime (kaplan2020scaling).
Empirically, learning curves exhibit an S-shaped transition when plotted in log–log coordinates (e.g. Figure 1). Motivated by this observation, we model the learning progress term $P(N,D)$ as a logistic function in $\log D$,
$P(N,D)=\frac{1}{1+\left(\frac{D}{D_{0}(N)}\right)^{\beta}},$ (18)
where $D_{0}(N)$ denotes the characteristic dataset scale at which half of the learnable capacity is realized, and is treated as a $N$-dependent parameter.
Combining the above components, we arrive at the following instantiated loss model:
$L(N,D)=L_{\infty}+\left(\frac{N_{0}}{N}\right)^{\alpha}+\frac{\lambda(N)}{1+\left(\frac{D}{D_{0}(N)}\right)^{\beta}}$ (19)
This is what we used to fit and extrapolation in Figure 15
D.2 Predictability and Extrapolation
(a) Intra-model prediction using the early 30% of training steps.
(b) Inter-model extrapolation for 72B under fixed shared shape.
Figure 15: Extrapolation of the loss decomposition model (Equation equation 19). (a) Intra-model prediction using partial learning-curve observations. (b) Inter-model extrapolation for 72B with global exponents fixed from smaller models.
We evaluate the extrapolation behavior of the loss–decomposition model (Equation equation 19) by applying it to the same experimental learning-curve data used in the main analysis, and by testing its performance under two complementary settings (Figure 15).
Intra-model prediction.
Using only the first 30% of training steps for each model, the fitted curves closely match the held-out portions(Figure 15a). This indicates that the internal S-shaped structure of the model provides a sufficiently strong inductive bias for completing a single learning curve from early observations. See Table 7 for detailed fitting result.
Inter-model extrapolation.
We fit all global exponents and the model-limited term using models up to 32B, and then extrapolate the resulting shared functional form to 72B by calibrating only two model-specific parameters, $\lambda(72\text{B})$ and $D_{0}(72\text{B})$. The extrapolated curve aligns well with the observed 72B trajectory across the full data range, reflecting that the functional shape inferred from smaller models remains compatible with larger-scale behavior under this light calibration.
The fit is reasonably strong, indicating that the proposed formulation may capture key structural tendencies of the underlying scaling behavior.
Table 7: Fitting details for $L(N,D)$ using only 30% of the training steps for each model.
Model Size Base
$D_{0}$ lambdalambda $R^{2}$
3B 5109.6316 0.0734 0.995
7B 3725.6374 0.1882 0.995
14B 4554.7619 0.2520 0.995
32B 3576.0323 0.3279 0.995
72B 5861.6228 0.3084 0.995
D.3 Discussion: Effective log–log slope
To relate the loss decomposition model to the slope-based formulation used in the main text, we examine the local behavior of $L(N,D)$ in log–log coordinates with respect to the data scale $D$. Specifically, we define the effective slope
$k(N,D)\;:=\;-\,\frac{\partial\log L(N,D)}{\partial\log D},$
which corresponds to the exponent in a local power-law approximation of the form $\log L\approx-k\log D+\text{const}$.
For the loss decomposition model (Equation equation 19), the induced $k(N,D)$ is a smooth function of $D$ that vanishes in both the low-data and high-data limits, and attains its maximum around the characteristic scale $D\approx D_{0}(N)$. Evaluating the slope at this point yields a natural definition of the maximal effective slope,
$k_{\max}(N)=\frac{K_{\max}}{1+S(N)},\qquad K_{\max}:=\frac{\beta}{2},\quad S(N):=\frac{2\!\left(L_{\infty}+(N_{0}/N)^{\alpha}\right)}{\lambda(N)}.$ (20)
The resulting $k_{\max}(N)$ depends only on the model size $N$ through the parameters of the loss decomposition, and is uniformly bounded by $K_{\max}=\beta/2$.
From this perspective, the slope function $k(N)$ adopted in the main analysis (Equation equation 2) can be interpreted as a parsimonious, low-parameter approximation to the effective maximal slope $k_{\max}(N)$ in Equation equation 20. This establishes structural consistency between the two descriptions: the loss decomposition model potentially captures the finer-grained $(N,D)$-dependent behavior, while the main-text formulation summarizes its dominant $N$-dependent trend in a compact form.
D.4 Conclusion
The loss decomposition model captures key empirical characteristics of RL post-training scaling behavior.
However, several components remain underdetermined, including the construction of the learnable capacity $\lambda(N)$, the dependence of the characteristic dataset scale $D_{0}(N)$ on model size, the role of the pretraining process, and the impact of mismatches between pretraining and post-training task settings.
For these reasons, we present this model as an appendix-level discussion rather than a core component of the main text. We encourage future work to build on this formulation to further investigate and refine scaling laws for LLM-based reinforcement learning.
Appendix E Response Length
Figure 16: Response length vs. Data size. Left: Base models. Right: Instruct models.